OpenStudy (anonymous):

You are cutting vegetables for a party. You have 1/4 lb of carrots, 2/7 lb of broccoli, and 5/6 lb of peppers. How many pounds of vegetables do you have in all?

OpenStudy (anonymous):

So basically you are adding all of the together. 1/4 + 2/7 + 5/6 Since they all have different denominators, you have to find a common one.

OpenStudy (anonymous):

What number will multiples of each of those numbers go into?

OpenStudy (anonymous):

would it the answer 1 1/4 thats what i just got!

OpenStudy (anonymous):

Do you know what you need to do? Or do you need a little more help?

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

Haha okay

OpenStudy (anonymous):

So we need to make all the denominators (the bottom numbers) the same. 4 * 21 = 84 7 * 12 = 84 6 * 14 = 84

OpenStudy (anonymous):

Now you need to multiply the numerators (the top numbers) times the same number you multiplied against the denominator. So 1*21, 2*12, 5*14... What do you get?

OpenStudy (anonymous):

You understand what you need to do?

OpenStudy (anonymous):

Yeah thank you so much! I understand now!!!

OpenStudy (anonymous):

Let me know what you get so I can help you if you didn't get it quite right... only if you want.

OpenStudy (anonymous):

HAHAHHA OKAY!

OpenStudy (anonymous):

hhaha okay so i got 1 1/4 again

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

Did you add them all together? the denominator should be 84

OpenStudy (anonymous):

WELL ITS NOT IN MY ASWENR CHOICES ANYWAYS

OpenStudy (anonymous):

If you added them all together you would get 115/84. That come out to 1 31/84, which in decimals would approx. equal 1.37

OpenStudy (anonymous):

What are your choices?

OpenStudy (anonymous):

1 1/4 LB, 5 LB, 2 1/2, 1 3/4

OpenStudy (anonymous):

That doesn't make any sense. Well 1 1/4 is the closest, so go with that.

OpenStudy (anonymous):

Okay then thank you =D

OpenStudy (anonymous):

viviansalem is that you in that pic? wow